OpenStudy (anonymous):

one number is four less than, five times another number. the total of the two numbers is 50. what are the numbers?

OpenStudy (anonymous):

x+5(x-4) = 50 x+5x-20 =50 6x-20=50 6x=30 x=5

OpenStudy (anonymous):

wouldn't 20 plus 50 be 70?